A man is six times as old as his son. Twenty years later, the man will be twice as old as his son. Find the age of the man when his son was born.

Answer by lwsshak3(11628)

A man is six times as old as his son. Twenty years later, the man will be twice as old as his son. Find the age of the man when his son was born.
***
let x=age of son
6x=age of father
..
20 yrs later:
x+20=age of son
6x+20=age of father
father twice as old as son
6x+20=2(x+20)
6x+20=2x+40
4x=20
x=5
6x=30
30-5=25
age of father when his son was born=25